OpenStudy (raden):

how to simply of : 1/(1+p+1/q) + 1/(1+q+1/r) + 1/(1+r+1/p)

OpenStudy (raden):

options : A. -1 B. 0 C. 1/(pq) D. -1 E. pq

OpenStudy (anonymous):

ur options are incorrect let r=1 ,p=1/2 and q=1 the answer will be 59/60

OpenStudy (raden):

sorry, pqr=1 so=?????

OpenStudy (anonymous):

well now put p=a/b and q=b/c and r=c/a answer will be 1

OpenStudy (raden):

ok i see.. thank u

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@RadEn come here

OpenStudy (raden):

ya, what is it @mukushla? oh iya mukushla, may i become ur friend in fb? because i want to invite u in my group "soul-mate-matika jilid 1". I think the group very" like have a member like u ^^